OpenStudy (anonymous):

factor completely: 10xr-3r-50x+15

OpenStudy (anonymous):

dang it this is confusing.

OpenStudy (anonymous):

not sure where the x and r go??

OpenStudy (anonymous):

10xr - 3r - 50x + 15 = 10xr - 50x + 15 - 3r = 10x(r - 5) + 3(5 - r) = 10x(r - 5) - 3(r - 5) = (10x - 3)(r - 5)

OpenStudy (anonymous):

(10x-3)(r-5) just it by trial and error

OpenStudy (anonymous):

will you please explain?

OpenStudy (anonymous):

step one and step two??

OpenStudy (anonymous):

how did you get 5??

OpenStudy (anonymous):

did you first switch

OpenStudy (anonymous):

10xr-50x-3r+15???????????

OpenStudy (anonymous):

then you asked what goes into 10 and 50?? same thing for 1 and 15????

OpenStudy (anonymous):

its the GCF 10 for 10 and 50????

OpenStudy (anonymous):

why 5 instead 10???????

OpenStudy (anonymous):

Just regroup it only Like for example you question ( 10xr-3r-50x+15) then (10xr-50x)-(3r-15) then common monomial of each term 10x(r-5)-3(r-5) then Copy the term outside the parenthesis and also the inside of the parenthesis (10x-3) (r-5)

OpenStudy (anonymous):

did u get it @Algebra2013

OpenStudy (anonymous):

I think I got it thank you very much!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

your welcome You just use the rule of grouping if u have 4 terms . then get it common monomial